### v1

# Fooling the Textual Fooler via Randomizing Latent Representations

Abstract
Despite outstanding performance in a variety of NLP tasks, recent studies have revealed that NLP models are vulnerable to adversarial attacks that slightly perturb the input to cause the models to misbehave. Among these attacks, adversarial word-level perturbations are well-studied and effective attack strategies. Since these attacks work in black-box settings, they do not require access to the model architecture or model parameters and thus can be detrimental to existing NLP applications. To perform an attack, the adversary queries the victim model many times to determine the most important words in an input text and to replace these words with their corresponding synonyms. In this work, we propose a lightweight and attack-agnostic defense whose main goal is to perplex the process of generating an adversarial example in these query-based black-box attacks; that is to fool the textual fooler. This defense, named AdvFooler, works by randomizing the latent representation of the input at inference time. Different from existing defenses, AdvFooler does not necessitate additional computational overhead during training nor relies on assumptions about the potential adversarial perturbation set while having a negligible impact on the model’s accuracy. Our theoretical and empirical analyses highlight the significance of robustness resulting from confusing the adversary via randomizing the latent space, as well as the impact of randomization on clean accuracy. Finally, we empirically demonstrate near state-of-the-art robustness of AdvFooler against representative adversarial word-level attacks on two benchmark datasets.
1 Introduction
In the last decade, deep neural networks have achieved impressive performance in the Natural Language Processing (NLP) domain. Several deep NLP models have reached state-of-the-art results in several NLP tasks (Devlin et al., 2019) using models such as Recurrent Neural Networks, Transformers, and Pretrained Language Models (PrLMs). However, several works (Iyyer et al., 2018) also reveal that deep NLP models can be fooled by the creation of adversarial examples (Jin et al., 2020). Adversarial examples are synthetically perturbed inputs that are optimized to increase the errors between the predictions of the model and the true labels while being imperceptible to human evaluators (Jin et al., 2020; Iyyer et al., 2018). These developments have sparked concerns about the security and robustness of deep neural networks deployed in NLP applications.
To generate adversarial examples, adversarial attack methods manipulate different aspects of the input sentence, from introducing character errors such as typos or visually similar characters (Gao et al., 2018a; Eger et al., 2019) and replacing words without significantly changing the original semantic in the perturbed inputs (Jin et al., 2020; Li et al., 2020) to recreating sentences with similar meanings using paraphrasing (Iyyer et al., 2018; Qi et al., 2021). Among these methods, adversarial-word substitutions (Alzantot et al., 2018) are some of the most widely studied and effective adversarial approaches. These methods query the victim model multiple times to find the corresponding adversarial perturbations of important words of an input text.
In response to the adversarial threat and to enhance the robustness of NLP models, numerous defenses against textual adversarial attacks have been developed (Zhu et al., 2020; Si et al., 2021; Wang et al., 2020; Ye et al., 2020; Shi et al., 2020b; Xu et al., 2020). The most popular approaches are adversarial training (Zhu et al., 2020; Dong et al., 2021; Madry et al., 2017), which augments the training data with adversarial examples using an additional optimization step, and randomized smoothing (Ye et al., 2020; Zeng et al., 2021), which replaces the model with its stochastic ensemble based on random perturbations of the input; both of these approaches require training modifications and additional non-trivial training computation (e.g., for adversarial training). Due to the discrete nature of texts, to create adversarial inputs, these defenses also substitute input words with adversarial words sampled from predefined perturbation sets constructed based on the potential attacks instead of gradient-based optimization. This substitution process also aims to preserve the original input semantics. Nonetheless, assuming knowledge of the potential attacks is often unrealistic and impractical.
In this paper, we propose a lightweight, attack-agnostic defensive method that can increase the robustness of NLP models against textual adversarial attacks. Our defense, called AdvFooler, randomizes the latent representation of the input at test time to fool the adversary throughout their attack, which typically involves iteratively sampling of discrete perturbations to generate an adversarial sample. Being a test-time defense with negligible computational overhead, AdvFooler also does not incur any training-time computation. Furthermore, as AdvFooler operates within the latent space of the model, it remains agnostic of the perturbation sets associated with potential attacks. While there are some randomization defenses in NLP (Ye et al., 2020; Zeng et al., 2021), they rely on randomized smoothing with necessary modifications to the model’s training to reduce the variance in the model’s outputs caused by randomizing the input or embedding space. The advantages of AdvFooler over the other defenses are shown in Table 1. Our contributions can be summarized as follows:
  • We propose a lightweight, attack-agnostic, and pluggable defensive method that hinders the attacker’s ability to optimize for adversarial perturbations leading to adversarial examples. Consequently, the attack success rate is significantly decreased, making the model more robust against adversarial attacks.
  • We provide important theoretical and empirical analyses showing the impact of randomizing latent representations on perplexing the attack process of the adversary.
  • We extensively evaluate AdvFooler through experiments on various benchmark datasets and representative attacks. The results demonstrate that AdvFooler is on par with state-of-the-art defenses while requiring few modeling assumptions, being pluggable, and incurring negligible additional computational overhead.
Table 1: Characteristics of all defenses. $\times$ or ✓ indicates the method lacks or has the specific characteristic, respectively. AdvFooler satisfies all the characteristics by adding lightweight randomization to the latent space.
Method no training required randomized trivial inference overhead no additional network pluggable ASCC $\times$ $\times$ $\times$ InfoBERT $\times$ $\times$ $\times$ FreeLB $\times$ $\times$ $\times$ TMD $\times$ $\times$ $\times$ $\times$ $\times$ SAFER $\times$ $\times$ $\times$ RanMASK $\times$ $\times$ $\times$ AdvFooler
2 Background
2.1 Adversarial attacks in NLP
Adversarial attacks in NLP can be divided into two categories, depending on what information the attacker has on the victim model. These categories are white-box attacks (Ebrahimi et al., 2018; Cheng et al., 2020) - those that require access to the model’s architecture and its parameters - and black-box attacks (Li et al., 2020; Jin et al., 2020; Li et al., 2018) - those that rely only querying the model for its output. Since the model user rarely shares the underlying model’s architecture of their NLP service, let alone their model’s parameters, white-box attacks have extremely limited practicality, compared to black-box attacks. In the paper, we will focus on the black-box, query-based attack setting, as it is a more realistic scenario in practice and is commonly studied in similar works (Zeng et al., 2021; Nguyen Minh & Luu, 2022; Zhang et al., 2022). Note that, there are also black-box, transfer-based attacks, which, however, still require knowledge of the model architecture for the attack to be effective (Yuan et al., 2021; Li et al., 2021a).
Black-box Query-based Attack. Given the text classification task, a model $f:\mathcal{R}^{d}\to\mathcal{R}^{|\mathcal{C}|}$ maps an input $x\in\mathcal{R}^{d}$ to a logit vector of dimension $|\mathcal{C}|$, where $\mathcal{C}$ is the set of label. The goal of textual adversarial attacks is to search for adversarial examples on which the model makes incorrect predictions. Specifically, given an input sentence $x$, a corresponding adversarial example $x^{\prime}$ can be crafted to satisfy the following objective:
$\arg\max_{c}f(x_{adv})\neq\arg\max_{c}f(x)$
[…]
where $d(x_{adv}-x)$ is the perceptual distance between $x$ and $x_{adv}$, and $\sigma$ is the maximum acceptable distance. The distance function $d$ can measure the number of perturbed words between the original and adversarial input (Gao et al., 2018b). Another popular distance is the semantic dissimilarity between a pair of texts (Iyyer et al., 2018; Li et al., 2020) measured as the cosine distance between the corresponding embedding vectors of the inputs extracted from Universal Sentence Encoder (USE) (Cer et al., 2018).
2.2 Adversarial word substitution
Among textual adversarial attacks, adversarial word substitution is one of the most widely studied approaches. These attacks create an adversarial example of an input text $x$ by first identifying the most important words in $x$ and then replacing those words with synonyms from their corresponding synonym sets to cause the victim model to misbehave while aiming to preserve the original semantic meaning. The synonym sets can be created using word embedding models (Mrkšić et al., 2016; Pennington et al., 2014) or using the predictions of a PrLMs given the input sentence (Li et al., 2020). One important feature of these attacks is that both of the aforementioned steps require querying the victim model many times to determine which words are important or which synonyms maximize the model’s prediction errors. In the first step, the adversary queries the model to calculate the importance score of each word in $x$ and select the most potential locations (with the highest importance scores) for later perturbations. To find the substituting synonyms that maximize the model’s prediction error, the attacks can either perform a greedy search (Li et al., 2020; Jin et al., 2020; Li et al., 2018) or combinatorial optimization algorithm (Zang et al., 2020; Alzantot et al., 2018). Similarly, this step involves a sequence of queries on the victim model. By introducing randomization to the latent representations of queries, our defensive mechanism disrupts the adversary’s ability to estimate important words, making their synonym-substituting process significantly harder during the attack.
2.3 Adversarial defense methods
To defend against textual adversarial attacks, several defenses (Wang et al., 2022) have been developed. The goal of an adversarial defense is to improve the robustness of the victim model; that is to achieve good performance on both clean and adversarial examples. Existing adversarial defenses can be divided into two categories: certified (Ye et al., 2020; Zeng et al., 2021; Jia et al., 2019) and empirical (Zhu et al., 2020; Wang et al., 2020; Dong et al., 2021; Li & Qiu, 2020; Zhou et al., 2021; Miyato et al., 2016). First introduced in Goodfellow et al. (2014), adversarial training is the most often studied empirical defense method. It provides additional regularization to the model and improves the model’s robustness when training it with adversarial samples. Several subsequent works study adversarial training for NLP tasks with adversarial examples created on the input space (Ren et al., 2019; Li et al., 2020; Jin et al., 2020). Some adversarial training works improve the model’s robustness by introducing adversarial perturbations in the embedding space (Zhu et al., 2020), or by incorporating inductive bias to prevent the model from learning spurious correlations (Wang et al., 2020; Madry et al., 2017). These methods effectively improve the model’s robustness without significant compromise to its clean accuracy.
In contrast to the empirical methods, certified defenses can provably guarantee model robustness even under sophisticated attackers. One popular certified-defense approach is randomized smoothing (Ye et al., 2020; Zeng et al., 2021), which constructs a set of stochastic ensembles from the input and leverages their statistical properties to provably guarantee robustness. In contrast, Interval Bound Propagation (IBP) methods (Shi et al., 2020a; Jia et al., 2019; Huang et al., 2019) utilizes axis-aligned bounds to confine adversarial examples. Despite their effectiveness and ability to certify robustness, both approaches rely on access to the synonym sets of potential attacks. Randomized-smoothing methods also modify the training phase to reduce the variance of the outputs of the ensemble. Due to the ensemble, they incur significant computational overhead in the inference phase. In contrast, our defense is lightweight and attack-agnostic, and can better control the variance of the outputs by randomizing the latent representations, instead of via randomization in the input space. A summary of the characteristics of AdvFooler and the existing defenses is shown in Table 1.
3 Methodology
3.1 Randomized latent-space defense against adversarial word substitution
Since the adversary relies on querying the victim model multiple times to find adversarial examples, receiving fluctuating feedback from the model will make it significantly harder for the adversary to find the optimal adversarial perturbations. To fool such an adversary, we introduce stochasticity to the model by randomizing the latent representation of an input $x$. Formally, for $h_{l}$ as the $l-$th layer of the model, we sample an independent noise vector $\epsilon$, which is added to $h_{l}(x)$ as input to the next layer of the model. Without loss of generality, $\epsilon$ is sampled from a Gaussian distribution $\mathcal{N}(0,\Sigma)$ with $\Sigma$ as a diagonal covariance matrix, or $\mathcal{N}(0,\nu I),\nu\in\mathcal{R}$. The detailed algorithm is presented in Algorithm 1.
Let $f_{\operatorname{AdvFooler}}$ be the proposed randomized model corresponding to the original $f$. When the variance of injected noise is low, we can assume that small noise slightly changes the output logits but does not shift the prediction of the model. In other words, the mean of the randomized model $f_{\operatorname{AdvFooler}}$ with input $x$ is exactly the prediction of $f$ for $x$. Consequently, adversarial samples of $f$ are adversarial samples of $f_{\operatorname{AdvFooler}}$.
Algorithm 1 AdvFooler’s Randomized Latent Defense
Input: a model $f$, an input sentence $x$, noise magnitude $\nu$.
      Output: output logit $l$.
1:$z_{0}=Emb(x)$
2:for  layer $l$ in $\{1...L\}$ of model $f$  do
3:     $\epsilon\sim\mathcal{N}(0,\nu I)$
4:     $z_{i+1}=h_{l}(z_{i}+\epsilon)$
5:end for
6:return $z_{L}$
3.2 Effect of randomizing latent representations on adversarial attacks
In textual adversarial attacks, the adversary first determines important words in a sentence $x=\{w_{0},\dots,w_{i}\}$ using the importance score, which is computed for each word $w_{i}$ as
$I_{w_{i}}=f_{y}(x)-f_{y}(x_{/w_{i}})$ (1)
where $f_{y}$ is the logit returned by the model $f$ w.r.t ground-truth label $y$. The adversary then selects words with the highest scores to perturb. Intuitively, this process emulates estimating the gradient of the importance score in the discrete space.
When introducing randomness to the latent space, the important scores are changed. Such changes mislead the attack to select a different set of words to perturb. Perturbing these “less important” words (those that have lower original importance scores before randomization) makes the synonym replacement phase significantly more challenging for the attack, thus reducing the attack’s success probability.
Theorem 1.
[…]
Theorem 1 (proof is in the supplementary material) states that randomly perturbing the hidden presentation leads to a randomized important score. When $\nu$ is high, the randomized importance scores have high variance, which can cause the attack to select the wrong important words.
3.3 Effect of randomizing latent representations on clean accuracy
If randomization induces substantial fluctuation in the model’s output, there is a possibility that the model might predict a wrong label, resulting in a decrease in clean accuracy. Randomized defenses such as SAFER (Ye et al., 2020) and RanMASK (Zeng et al., 2021) introduce randomness to the model in the input space (i.e., randomizing the input tokens). While they are also effective against textual adversarial attacks, the input space of NLP models is discrete. Perturbing discrete inputs can lead to significant changes in the model’s outputs (a fact that we will empirically prove in the next section), which results in a different model’s prediction. To mitigate this problem, existing randomized input defenses rely on randomized smoothing: train the model with perturbed input and apply ensemble at inference to reduce the variance in the model’s output. Nevertheless, these approaches are computationally expensive and require access to the training phase. Conversely, by randomizing the latent representations, AdvFooler induces a significantly smaller variance in the model’s output. Furthermore, AdvFooler gives the defender flexibility to select the suitable induced variance without requiring access to the training process. Specifically, given a pretrained classification model and a small clean test set, the defender can select the noise scale $\nu$ at which the variance causes the clean accuracy drops by a chosen percentage (e.g., in our experiments, it is 1%).
3.4 Empirical analysis
In this section, we empirically study the effect on the model’s performance and attack process of various randomized defenses, including AdvFooler.
Empirical effect on model’s performance. We compute the differences in cross-entropy losses between those of the randomized model and those of the base model for each sample in the AGNEWS test dataset, which expresses how the predictions vary, under different randomization approaches. Figure 2 shows the loss changes in SAFER, RanMASK, and AdvFooler. As we can observe, in SAFER and RanMASK, perturbing the input tokens induces a higher variance in the loss compared to AdvFooler. Consequently, both of these methods lead to significant drops in the model’s performance without adversarial training and ensemble.
Empirical effect on fooling the attacks. To demonstrate that AdvFooler can fool the adversary into selecting non-important words, we calculate the important score of each word for multiple samples from the AGNEWS dataset. The important words before and after applying AdvFooler are presented in Figure 2. As we can observe, AdvFooler causes the attack to select different important words, many of which are originally unimportant. Specifically, in the first sentence, the token “Airlines” is the most important. After adding random noise to the latent representation, AdvFooler changes the important score for each token enough to mislead the adversary into thinking a different token (“Yesterday” token) is the most important token. A similar phenomenon can be seen in the second sentence, where the important scores change significantly, and unimportant tokens become important ones after randomization. Note that, these experiments use the random noise scale $\nu$ that induces at most 1% clean performance drop.
Figure 1: Loss changes when randomizing input (RanMASK/SAFER) and latent space (AdvFooler).
Figure 1: Loss changes when randomizing input (RanMASK/SAFER) and latent space (AdvFooler).
Figure 2: Illustration of each word’s important score when calculated with and without AdvFooler.
4 Experiments
4.1 Experimental setup
Table 2: The robustness performance ofAdvFooler and other defenses on AGNEWS. The best and second-best performances are bolded and underlined, respectively.
Models Clean(Drop%) TextBugger TextFooler BERT-Attack
AuA(ASR%) #Query AuA(ASR%) #Query AuA(ASR%) #Query
BERT-base (baseline) 94.20 35.20 (62.63) 351.92 36.80 (60.93) 317.27 46.50 (50.64) 337.84
+ASCC 90.07 (--4.13) 39.30 (55.94) 284.58 41.20 (53.81) 262.07 45.90 (48.54) 263.59
+FreeLB 95.07 ($+$0.87) 47.30 (49.89) 394.19 49.90 (47.14) 356.69 52.00 (44.92) 370.82
+InfoBERT 95.01 ($+$0.81) 45.20 (52.07) 373.83 47.50 (49.63) 336.77 56.00 (40.62) 366.40
+TMD 94.36 ($+$0.16) 47.80 (49.04) 833.43 51.60 (44.81) 744.87 52.60 (43.74) 766.72
+RanMASK 90.14 (--4.06) 52.50 (41.60) 582.21 54.60 (38.93) 511.98 61.10 (31.73) 595.57
+SAFER 94.42 ($+$0.22) 43.30 (53.74) 370.77 46.70 (50.05) 333.17 51.40 (45.03) 357.47
+AdvFooler (ours) 93.67 (--0.53) 50.10 (45.90) 819.44 50.10 (45.90) 701.05 53.40 (42.89) 752.61
RoBERTa-base (baseline) 95.05 38.40 (58.93) 371.20 41.00 (56.15) 335.88 47.30 (49.41) 372.99
+ASCC 92.61 (--2.44) 49.50 (46.37) 372.91 51.00 (44.75) 330.69 55.70 (39.65) 374.58
+FreeLB 95.03 (--0.02) 50.50 (46.22) 448.97 52.00 (44.62) 402.14 58.20 (38.02) 442.97
+InfoBERT 94.97 (--0.08) 47.00 (50.11) 394.56 48.90 (48.09) 357.44 55.00 (41.61) 397.35
+TMD 93.70 (--1.35) 52.80 (43.59) 878.22 55.20 (41.09) 766.74 57.90 (38.14) 805.19
+RanMASK 90.41 (--4.64) 52.30 (41.50) 575.94 54.20 (39.37) 516.98 60.00 (32.43) 594.93
+SAFER 94.67 (--0.38) 47.70 (49.15) 377.67 50.30 (46.43) 337.81 53.20 (43.22) 357.56
+AdvFooler (ours) 94.21 (--0.84) 51.70 (44.88) 804.99 53.30 (42.81) 717.31 57.80 (38.38) 767.47
Evaluation Metrics. We evaluate the performance of AdvFooler using four evaluation metrics: Clean accuracy (Clean%), Accuracy under attack (AuA%), Attack success rate (ASR%), Average number of queries (#Query). Cleanmeasures the impact of a defense method on the clean performance of the model. Both AuAand ASRmeasure the robustness of the model under attack. #Query measures how many times, on average, the adversary has to query the model to find adversarial examples; the higher #Query is, the more difficult the attack is.
Datasets. We evaluate AdvFooler on two benchmark datasets: AG-News Corpus (AGNEWS(Zhang et al., 2015) and Internet Movie Database (IMDB(Maas et al., 2011). AGNEWS is a classification dataset of news articles created from the AG’s corpus [^1] http://groups.di.unipi.it/~gulli/AG_corpus_of_news_articles.html. It contains 120,000 training samples and 7,600 test samples where a sample belongs to one of four classes: World, Sports, Business, and Sci/Tech. IMDB is a binary sentiment classification dataset of reviews extracted from the IMDB website. This dataset contains 50,000 samples and is split equally into the training and test sets.
Model Architecture. To assess the generality of AdvFooler on model architectures, we evaluate it on two state-of-the-art, pre-trained language models: BERTbase (Devlin et al., 2019) and RoBERTabase (Liu et al., 2019). BERT is a Transformer-based language model that outperforms other models in various benchmarks at the time of its release and is widely used for text classification tasks. RoBERTa is a variant of BERT with an improved training process and performance.
Adversarial Attacks and Parameters. We choose the following widely used and state-of-the-art adversarial attacks to evaluate our defense: TextFooler (Jin et al., 2020), TextBugger (Li et al., 2018), and BERT-Attack (Li et al., 2020). TextFooler utilizes the word embeddings from (Mrkšić et al., 2016) to generate the set of synonyms for each word, uses both part-of-speech and semantic similarity, and selects words for replacements using important scores. BERT-Attack is similar to TextFooler but uses a pretrained BERT to generate the set of synonyms for a word, given the context of the sentence. TextBugger locates the tokens for perturbations by calculating the important scores for each sentence in the input and then calculating the important score of each token in the sentence. In addition to a synonym set for word substitution, TextBugger also introduces word augmentation - perturbing the characters in words. For a fair evaluation, the adversarial attacks used in the evaluation will be based on Li et al. (2021b), which is commonly used by other papers for benchmarking adversarial defenses (Nguyen Minh & Luu, 2022; Zhang et al., 2022)
Table 3: The robustness performance ofAdvFooler and other defenses on IMDB. The best and second-best performances are bolded and underlined, respectively.
Models Clean(Drop%) TextBugger TextFooler BERT-Attack
AuA(ASR%) #Query AuA(ASR%) #Query AuA(ASR%) #Query
BERT-base (baseline) 92.14 9.20 (90.08) 500.47 11.90 (87.16) 439.15 8.90 (90.40) 366.52
+ASCC 88.48 (--3.66) 13.00 (85.68) 597.07 16.90 (81.20) 529.55 7.70 (91.52) 416.50
+FreeLB 92.33 ($+$0.19) 25.80 (71.74) 776.31 28.90 (68.35) 670.20 21.80 (76.12) 549.85
+InfoBERT 91.71 (--0.43) 22.50 (75.52) 719.98 25.30 (72.47) 645.03 20.90 (77.26) 510.46
+TMD 92.14 (--0.00) 40.40 (56.47) 3251.52 45.10 (51.35) 2735.58 36.20 (60.95) 2464.83
+RanMASK 92.61 ($+$0.47) 31.00 (66.38) 2740.39 35.80 (61.13) 2392.46 33.10 (64.33) 2463.24
+SAFER 92.12 (--0.02) 46.10 (50.00) 1455.25 50.50 (45.64) 1262.35 43.10 (52.74) 1133.61
+AdvFooler (ours) 91.90 (--0.24) 42.40 (53.41) 3261.41 49.10 (47.32) 2759.37 40.70 (55.76) 2645.36
RoBERTa-base (baseline) 93.23 6.90 (92.63) 517.58 11.40 (87.79) 456.35 8.20 (91.29) 439.79
+ASCC 92.62 (--0.61) 14.70 (84.31) 770.35 20.20 (77.92) 606.14 15.20 (83.57) 548.34
+FreeLB 94.20 ($+$0.97) 25.40 (72.95) 887.31 29.80 (68.30) 726.84 23.10 (75.19) 647.98
+InfoBERT 94.18 ($+$0.95) 20.90 (78.14) 684.64 27.60 (70.61) 583.97 15.50 (83.53) 518.17
+TMD 93.22 (--0.01) 66.10 (29.30) 4799.56 66.40 (28.83) 3477.32 56.00 (40.36) 3330.08
+RanMASK 94.33 ($+$1.10) 49.40 (47.45) 3611.03 54.10 (43.29) 2951.45 44.20 (53.67) 2706.83
+SAFER 93.75 ($+$0.52) 62.30 (34.21) 2059.37 63.60 (32.27) 1874.79 54.10 (41.77) 1415.56
+AdvFooler (ours) 92.69 (--0.54) 62.20 (33.55) 4205.77 63.70 (32.02) 3255.35 49.40 (47.00) 3141.88
Baseline Defenses. We compare AdvFooler against various types of defense methods. For empirical defenses, we select Adversarial Sparse Convex Combination (ASCC) (Dong et al., 2021), InfoBERT (Wang et al., 2020), FreeLB (Zhu et al., 2020), and Textual Manifold Defense (TMD) (Nguyen Minh & Luu, 2022). ASCC, InfoBERT, and FreeLB improve the model’s robustness either by employing adversarial training or introducing additional regularization to the model’s training process. ASCC models the word substitution attack space as a convex hull and leverages a regularization term to generate adversarial examples. Different from ASCC, FreeLB introduces adversarial perturbations to word embeddings and minimizes the corresponding adversarial loss around input samples. InfoBERT adds two regularization terms to eliminate noisy information between the input and hidden features and to increase the correlations between the local and global features. Textual Manifold Defense introduces a defense mechanism that trains a mapping function to project the embeddings of input sentences to a set of predefined manifolds to reduce the effects of the adversarial examples. For certified defenses, RanMASK (Zeng et al., 2021) and SAFER (Ye et al., 2020) are selected. Given an input, both RanMASK and SAFER construct a set of randomly perturbed inputs, using synonym substitution for SAFER and token masking for RanMASK, respectively. Then, they leverage the statistical properties of the predicted output to achieve better model robustness.
Implementation Details. We follow the implementation guidelines for evaluating adversarial attacks and defenses in Li et al. (2021b): (1) the maximum percentage of modified words $\rho_{max}$ for AGNEWS and IMDB are 0.3 and 0.1, respectively, (2) the maximum number of candidate replacement words $K_{max}$ is set to 50, (3) the maximum percentage of modified words $\rho_{max}$ for AGNEWS and IMDB must be 0.3 and 0.1, respectively, and (4) the maximum number of queries to the victim model is $Q_{max}=K_{max}*L$, where $L$ is the length of, or the number of tokens in, the input sentence. Cleanis calculated on the entire test set from each dataset, while the robustness metrics AuA%, ASR%, and #Query are computed on 1,000 randomly chosen samples from the test set. All attacks in the experiments are based on the implementation of TextAttack (Morris et al., 2020).
4.2 Defense performance against textual adversarial attacks
We report the defense performance on AGNEWS in Table 2. As we can observe in this table, on the base model BERT, AdvFooler outperforms all adversarial training techniques (ASCC, FreeLB, and InfoBERT) and the randomized smoothing method, SAFER. The robustness of AdvFooler is comparable to TMD’s robustness and slightly lower than the robustness of another randomized-smoothing defense, RanMASK. However, the clean accuracy in RanMASK drops significantly, approximately 4%, while the clean accuracy in AdvFooler is guaranteed to be within 1% of the base model. Note that, all of the evaluated textual defenses require access to the training data. On the base model RoBERTa, AdvFooler achieves similar robustness to the best defenses, but without the additional computational overhead (see supplementary material for the computational overhead experiment) or access to the training data.
For IMDB’s results in Table 3, we can observe similar robustness results. AdvFooler is effective against the attacks and achieves comparable results to the best defenses (SAFER and TMD).
Difficulty of Adversarial Attacks against AdvFooler. For all the considered adversarial attacks, the number of queries to locate important words (the first step) of an input text is the same with or without the defenses. Thus, the higher number of queries in an experiment is a result of a more difficult synonym-replacement phase. This also explains considerably smaller average numbers of queries for the baseline (i.e., without any defense) in Tables 2 and 3, compared to the protected models.
We can observe, in Tables 2 and 3, that the average numbers of queries for the adversarial attacks to be successful against AdvFooler are significantly higher than those in all the defenses except TMD in a few cases. The results demonstrate the effectiveness of fooling the attacks in AdvFooler. As explained in Section 3, randomizing the important scores just enough makes the adversary select a different set of important words, which leads to a more challenging synonym-replacement process with a significantly higher number of queries.
AdvFooler with other attacks. In the supplementary materials, we provide additional experiments evaluating AdvFooler  against other types of attacks, including black-box hard-label (Appendix LABEL:sec:hard-label) and white-box attacks (Appendix LABEL:sec:white-box), even though they do not align with the threat model studied in this work and several related works (Nguyen Minh & Luu, 2022; Zhang et al., 2022; Zeng et al., 2021; Li et al., 2023) to AdvFooler. The results show that AdvFooler  is still effective in defending against these attacks, further highlighting the versatility of AdvFooler.
4.3 Robustness from randomizing different latent spaces
Figure 3: The robustness of AdvFooler when randomizing different layers of the model on AGNEWS.
In this section, we study the model’s robustness when randomizing different layers of the model. We also decided to study the effects of adding noise to only one layer of the model. For comparison, Specifically, we randomize only the first, middle, and last layers and all layers of the model. The model’s robustness against TextFooler is reported in Figure 3. As we can observe from this figure, randomizing the early layer of the model is more effective than randomizing the later layers of the model. More importantly, adding noise to all the layers leads to a significantly higher accuracy under attack and a higher average number of queries.
4.4 Trade-off between clean accuracy and robustness
((a)) AGNEWS
((b)) IMDB
Figure 4: Clean Accuracy and Accuracy under Attack (AuA) when using different noise scales $\nu$.
As discussed in Section 3, our approach allows the defender to flexibly trade-off between the clean accuracy (or variance in the output logits) and the model’s robustness, a feature that is challenging to perform in other defenses. In AdvFooler, this is accomplished via tuning the noise scale $\nu$. A large $\nu$ value leads to both a large variance of the model’s output, which can lead to a lower clean accuracy, and a high chance of fooling the adversarial attacks.
We report the model’s clean accuracy and robustness when varying the value of $\nu$ in Figure 4. As we can observe, increasing $\nu$ generally increases the model’s robustness against all the evaluated attacks while the clean accuracy only slightly changes. However, when $\nu$ reaches a certain value, the clean accuracy begins to decrease, at which point the robustness of the model also decreases. The plot also shows that the noise scale when both accuracy under attack and clean accuracy start to decrease would also be the noise scale that makes the model more robust to all types of attack. This suggests that, in practice, the defender selects as large $\nu$ value as possible for a certain small drop of clean accuracy. For example, in our experiments, we select the $\nu$ value at which the clean accuracy drops by at most 1% using the test set.
Table 4: Clean Accuracy and Accuracy under Attack of TextBugger (TB), TextFooler (TF), and BERT-Attack (BA) for FreeLB- and InfoBERT-trained models, with and without using AdvFooler .
Dataset Models Clean (%) AuA(%)
TB TF BA
AGNEWS FreeLB 95.07 47.3 49.9 52.0
FreeLB+AdvFooler 94.61 62.5 63.2 63.2
InfoBERT 95.01 45.2 47.5 56.0
InfoBERT+AdvFooler 94.34 64.3 64.8 67.1
IMDB FreeLB 92.33 25.8 28.9 21.8
FreeLB+AdvFooler 92.22 49.6 50.3 40.7
InfoBERT 91.71 22.5 25.3 20.9
InfoBERT+AdvFooler 91.46 46.6 51.7 42.1
4.5 AdvFooler with adversarially trained models
As discussed previously, AdvFooler can be used to improve the robustness of any existing pretrained model; i.e., AdvFooler is a pluggable defense. In fact, AdvFooler can also be combined with other types of defenses to provide another layer of protection and further improve the model’s robustness. In this section, we study the effectiveness of AdvFooler in conjunction with adversarial training methods, including FreeLB and InfoBERT.
Table 4 shows the defense performance before and after applying AdvFooler on adversarially trained models using FreeLB and InfoBERT. We can observe that AdvFooler significantly improves the robustness of the adversarially trained models, with up to 20% improved accuracy under attack in some experiments. While AdvFooler can also be combined with other defensive mechanisms, we leave these to future works.
5 Conclusion
In this work, we proposed a lightweight, attack-agnostic defense, AdvFooler, that can improve the robustness of NLP models against textual adversarial attacks. Different from existing defenses, AdvFooler does not incur any training overhead nor relies on assumptions of the potential attacks. The main idea of AdvFooler is to mislead word-level perturbation-based adversary into selecting unimportant words by randomizing the latent space of the model, resulting in a significantly more challenging synonym-replacement process. We then provide both theoretical and empirical analyses to explain how AdvFooler fools the adversary, as well as its effect on the model’s accuracy. Our extensive experiments validate that AdvFooler improves the robustness of NLP models against multiple state-of-the-art textual adversarial attacks on two datasets. Finally, being a pluggable defense, AdvFooler can also be combined with existing defenses, such as adversarial training, to further protect the NLP models against these threats.